OpenStudy (anonymous):

So I have a small question about inconsistent equations. So I solved a problem using elimination by addition and turned out with this order pair (4,3) and I had to check it with these two equations 4x-2y=10 6x-3y=-3 I know the pairs work with the first equation but it does not work with the second equation. Does this make this equation inconsistent? Thanks!

OpenStudy (anonymous):

Do you anything about matrix algebra?

OpenStudy (anonymous):

Nope..

OpenStudy (anonymous):

Alright, here is an elementary proof why this two equation is inconsistent. \( 4x-2y=10 \implies 2x-y=5 \) \( 6x-3y=-3 \implies 2x-y=-1 \) This means that the two equation represents two parallel lines and hence they will never intersect so no solutions.

OpenStudy (anonymous):

Ah, I see. Thank you!

OpenStudy (anonymous):

Would it possible for you to give me an example of an inconsistent equation?

OpenStudy (anonymous):

\( 2x-y=5 \) \( 4x-2y=4\)

OpenStudy (anonymous):

What about a dependent equation? Let's say I have an equation that cancels everything out. -4x-6y=-12 4x+6y=12 Would that be dependent?